The total area of four sides of a rectangular box of height 3 cm is 72 cm^2.If the length were to be increased by 10% and the breadth decreased by 20%, this area would remain unchanged.Find the length and breadth of the box.
Answer by ankor@dixie-net.com(22740)

The total area of four sides of a rectangular box of height 3 cm is 72 cm^2.
therefore
2(3L) + 2(3W) = 72
simplify, divide by 2
3L + 3W = 36
:
If the length were to be increased by 10% and the breadth decreased by 20%, this area would remain unchanged.Find the length and breadth of the box.
A 10% increase in Length would be 1.1L
A 20% decrease in width would be .8W
Therefore
2(3*1.1L) + 2(3*.8W) = 72
simplify you have
3.3L + 2.4W = 36
Multiply the first equation by .8, subtract it from the above equation
3.3L + 2.4W = 36
2.4L + 2.4W = 28.8
-------------------
.9L = 7.2
L = 7.2/.9
L = 8 cm, the original length
then
3(8) + 3W = 36
3W = 36-24
W = 12/3
W = 4 cm is the original width
:
:
I'll let you confirm this solution in the 2nd statement
1.1(8) = 8.8 cm increase length
.8(4) = 3.2 cm the decreased width